# Joint multi-band deconvolution for Euclid and Vera C. Rubin images

### v1

@@
Joint multi-band deconvolution for Euclid and Vera C. Rubin images
@@
With the advent of surveys like Euclid and Vera C. Rubin, astrophysicists will have access to both deep, high-resolution images, and multi-band images. However, these two conditions are not simultaneously available in any single dataset. It is therefore vital to devise image deconvolution algorithms that exploit the best of the two worlds and that can jointly analyze datasets spanning a range of resolutions and wavelengths. In this work, we introduce a novel multi-band deconvolution technique aimed at improving the resolution of ground-based astronomical images by leveraging higher-resolution space-based observations. The method capitalizes on the fortunate fact that the Vera C. Rubin $r$-, $i$-, and $z$-bands lie within the Euclid $VIS$ band. The algorithm jointly deconvolves all the data to turn the $r$-, $i$-, and $z$-band Vera C. Rubin images to the resolution of Euclid by enabling us to leverage the correlations between the different bands. We also investigate the performance of deep learning-based denoising with DRUNet to further improve the results. We illustrate the effectiveness of our method in terms of resolution and morphology recovery, flux preservation, and generalization to different noise levels. This approach extends beyond the specific Euclid-Rubin combination, offering a versatile solution to improve the resolution of ground-based images in multiple photometric bands by jointly using any space-based images with overlapping filters.
@@
Deconvolution – EuclidVera C. Rubin – HST
@@ 1 Introduction
High spatial resolution, high signal-to-noise ratio, and broad wavelength coverage are all essential to most observational challenges in astrophysics. However, it is difficult, or even impossible, to make these three miracles happen simultaneously. Space telescopes, although free of atmospheric turbulence, are limited in size. Ground-based telescopes can deliver high signal-to-noise data but are affected by atmospheric turbulence and have a higher sky background. In addition, blurring by the instrumental or atmospheric Point Spread Function (PSF) differs in each type of data and varies from band to band. To capitalise on the strengths of all types of telescopes and data, it is crucial to develop robust deconvolution techniques that can remove blurring by the PSF and that can optimize the signal-to-noise ratio of the final reconstruction, by combining all observations accounting for the different bands and PSFs.
@@
Due to the presence of noise, image deconvolution is a challenging ill-posed inverse problem that requires regularization for a well-defined solution. Early approaches in the field acknowledged this issue, proposing solutions such as minimising the Tikhonov function (Tikhonov & Arsenin, 1977) or maximising the entropy of the solution (Skilling & Bryan, 1984). Bayesian methods also emerged, including the Richardson-Lucy algorithm applied to early Hubble Space Telescope (HST) data (Richardson, 1972; Lucy, 1974). A novel approach proposed by Magain et al. (MCS; 1998) separated point sources from extended ones and used a narrow PSF for deconvolution to achieve an improved resolution suitable for the chosen pixel sampling. This approach was enhanced with wavelet regularization for the extended channel (Cantale et al., 2016), and further refined by Michalewicz et al. (STARRED; 2023), who employed Starlets, an isotropic wavelet basis (Starck et al., 2015), to regularize the solution. There have also been efforts to jointly deconvolve multiple astronomical observations of the same sky region (Donath et al., 2023). Furthermore, Ingaramo et al. (2014) explored the combination of multiple sources by demonstrating the application of Richardson-Lucy deconvolution to merge high-resolution, high-noise images with low-resolution, low-noise images.
@@
A notable advancement in astronomical deconvolution was the use of deep learning. Once trained, deep learning-based methods also offer significant computational efficiency compared to traditional approaches. U-Nets (Ronneberger et al., 2015) have gained popularity for their nonlinear processing capabilities and multi-scale architecture. Expanding on U-Nets, Sureau et al. (2020) introduced the Tikhonet method for deconvolving optical galaxy images, demonstrating superior performance over sparse regularization methods in terms of mean squared error (MSE) and a shape criterion that assesses galaxy ellipticity. Nammour et al. (2022) enhanced Tikhonet by incorporating a shape constraint into the loss function, achieving further improvements. Another powerful architecture, Learnlet (Ramzi et al., 2023), combines the strengths of wavelets and U-Nets while offering a fully interpretable neural network with minimal hallucination. In our previous work (Akhaury et al., 2022, 2024), we proposed a two-step deconvolution framework and investigated the performance of CNN and transformer-based denoisers. We concluded that a Swin transformer-based U-Net (SUNet, Fan et al. (2022)) outperformed a CNN-based U-Net (Ronneberger et al., 2015) in terms of normalised mean sqaured error(NMSE) and structural similarity index measure (SSIM).
@@
While deconvolution is primarily used to reconstruct galaxy images at high spatial resolution in each photometric band independently, there are scenarios, particularly at low signal-to-noise ratios, where joint multi-band deconvolution can enhance the detection and characterization of systems. One such potential application is the joint multi-band deconvolution of Euclid and Vera C. Rubin images. The Vera C. Rubin Observatory is set to deliver a dataset of $500$ petabytes across multiple optical frequency bands, while Euclid will observe images spanning the optical and infrared spectrum. Interestingly, the Euclid $VIS$-band (central frequency $=715$ nm) overlaps with three of the Vera C. Rubin filters: $r$, $i$, $z$. As a space-based satellite, Euclid will produce images with sharper details due to its narrower PSF compared to Vera C. Rubin. A related study by Joseph et al. (2021) also involves jointly modelling simulated images that model observations from both Euclid and Vera C. Rubin.
@@
In this work, we present a novel multi-band deconvolution technique designed to enhance the resolution of ground-based astronomical images by leveraging higher-resolution space-based observations. Our approach, which focusses on the joint deconvolution of Vera C. Rubin and Euclid images, effectively exploits the overlapping spectral coverage of the Rubin $r$-, $i$-, and $z$-bands with the Euclid $VIS$-band. By utilizing the Euclid $VIS$-band image as a term that provides additional information, our technique ensures that the deconvolved Rubin images retain high spatial resolution and accurate photometric measurements. The integration of deep learning-based denoising further enhances the quality of the deconvolved outputs, reducing background noise without altering the main structures of the galaxies. We generate the realistic Euclid and Vera C. Rubin simulations from HST cutouts of varying magnitudes extracted from the GOODS-N and GOODS-S surveys (Retzlaff, J. et al., 2010). The simulated Euclid-like $VIS$-band PSF was obtained from Liaudat et al. (2022), and the simulated Rubin-like $r$-, $i$-, and $z$-band PSFs were obtained from Abolfathi et al. (2021). Our method demonstrates effectiveness in terms of resolution recovery, flux preservation, and generalization across different noise levels. Through our joint deconvolution approach, we achieve resolution recovery in Vera C. Rubin simulated images close to that of HST, a feat nearly impossible with independent deconvolution of each photometric band. The potential applications of our method go beyond the Euclid-Rubin pair, providing a flexible solution to enhance the resolution of ground-based images in multiple photometric bands with any overlapping space-based filter band. This versatility is especially important as large-scale astronomical surveys gather increasing amounts of data, creating a need for effective and reliable deconvolution techniques.
@@
In Section 2, we describe the deconvolution problem and introduce our proposed solution. The methodology for generating our dataset is detailed in Section 3. We then present the outcomes of our deconvolution algorithm in Section 4. Finally, Section 5 presents our conclusions. To support reproducible research, the codes utilised in this article are publicly accessible in Section 6.
@@ 2.1 The forward model
For the three Vera C. Rubin filters, let $\mathbf{y}_{r},\mathbf{y}_{i},\mathbf{y}_{z}\in\mathbb{R}^{n\times n}$ be the corresponding observed images and $\mathbf{h}_{r},\mathbf{h}_{i},\mathbf{h}_{z}\in\mathbb{R}^{n\times n}$ be the PSFs. If $\mathbf{x}_{r}^{t},\mathbf{x}_{i}^{t},\mathbf{x}_{z}^{t}\in\mathbb{R}^{n\times n}$ denote the corresponding target images, $\ast$ denotes the convolution operation, and $\eta_{r},\eta_{i},\eta_{z}\in\mathbb{R}^{n\times n}$ denote additive noise, the observed Rubin images can then be modelled as:
@@
$\mathbf{y}_{z}$ $=\mathbf{h}_{z}\ast\mathbf{x}_{z}^{t}+\eta_{z}$ (3)
As for Euclid, let $\mathbf{y}_{euc}\in\mathbb{R}^{n\times n}$ be the observed image, $\mathbf{x}^{t}_{euc}\in\mathbb{R}^{n\times n}$ be the target image, and $\mathbf{h}_{euc}\in\mathbb{R}^{n\times n}$ be the PSF. If $\eta_{euc}\in\mathbb{R}^{n\times n}$ denotes additive noise and $\alpha_{r},\alpha_{i},\alpha_{z}\in\mathbb{R}^{n}$ denote the corresponding fractional flux contribution from each Vera C. Rubin filter, the target and the observed images can be modelled as:
@@
$\mathbf{y}_{euc}$ $=\mathbf{h}_{euc}\ast\mathbf{x}_{euc}^{t}+\eta_{euc}$ (5)
The motivation behind taking the weighted sum of the Vera C. Rubin images to model the Euclid image can be seen in Figure 1, which shows the overlap between the Rubin and Euclid filters.
Figure 1: The filter curves for Euclid and Vera C. Rubin, which display the relative filter transmission as a function of the wavelength. The Euclid $VIS$-band overlaps with three of the Vera C. Rubin filters: $r$, $i$, $z$.
@@ 2.2 The proposed solution
$L_{z}(\mathbf{x}_{z})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{z}\ast\mathbf{x}_{z}-\mathbf{y}_{z}}{\sigma_{z}}\right\|_{F}^{2}+\lambda_{z}\textbf{K}$ (8)
where $\text{where }\textbf{K}=\left\|\frac{\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}}{\sigma_{euc}}\right\|_{F}^{2}$ (9)
The first terms in equations 6-8 represent the data fidelity terms for each respective band, with $\sigma_{r}$, $\sigma_{i}$, $\sigma_{z}$ being the corresponding noisemaps. The second terms are the constraining terms that enforce the condition that the sum of images in the Rubin $r$-, $i$-, and $z$-band images equals the $Euclid$ $VIS$-band images, as expected from Figure 1 after flux calibration. Within these constraining terms, the individual images $x_{r}$, $x_{i}$, and $x_{z}$ are multiplied by their respective fractional flux contributions $\alpha_{r}$, $\alpha_{i}$, and $\alpha_{z}$, which represent the fractional area overlaps between their corresponding filter curves. The values of $\alpha_{r}$, $\alpha_{i}$, and $\alpha_{z}$ are obtained by integrating the area under the curves in Figure 1 and normalising them to sum up to one. The resulting values are $\alpha_{r}=0.3785$, $\alpha_{i}=0.3468$, and $\alpha_{z}=0.2746$. The denominator, $\sigma_{euc}$, denotes the $Euclid$ image noisemap. The choice of the multiplicative hyperparameters $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ is described in Section 2.3.
@@ 2.3 Hyper-parameter tuning
The hyper-parameters $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ are determined by varying the ratios between the second and first terms in equations 6-8. These ratios indicate the contribution coming from the constraining term. We varied the ratios from 0 (no contribution from the constraining term) to 1 (equal contribution from the constraining term) in steps of 0.01. From this experiment, the optimal solution yields the lowest MSE with a ratio of 0.3 for all three photometric bands. Subsequently, the values of $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ are computed from these ratios by dividing them by K (equation 9).
@@ 2.4 Optimization
The aim is to find optimal solutions $\mathbf{\hat{x}}_{r}$, $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{x}}_{z}$ that minimise the individual loss functions:
$\mathbf{\hat{x}}_{\{r,i,z\}}=\underset{\mathbf{x}_{\{r,i,z\}}}{\operatorname{argmin}}\hskip 2.0ptL_{\{r,i,z\}}(\mathbf{x}_{\{r,i,z\}})$
@@
$\mathbf{x}_{\{r,i,z\}}^{[k+1]}=\mathbf{x}_{\{r,i,z\}}^{[k]}-\beta_{\{r,i,z\}}\nabla L_{\{r,i,z\}}(\mathbf{x}_{\{r,i,z\}}^{[k]})$ (10)
where $\mathbf{x}^{[k]}$ denotes the variable $\mathbf{x}$ at $k^{th}$ iteration, and $\beta_{r},\beta_{i},\beta_{z}\in\mathbb{R}^{n}$ are the step-sizes chosen such that convergence is guaranteed (described in more detail in Section 2.5). While computing equation 10 for one band, it is assumed that the other two bands are known and remain constant. The gradients of the loss functions 6-8 are given by:
@@
$\nabla L_{z}(\mathbf{x}_{z})$ $=\frac{\mathbf{h}_{z}^{\top}\ast(\mathbf{h}_{z}\ast\mathbf{x}_{z}-\mathbf{y}_{z})}{\left\|\mathbf{\sigma}_{z}\right\|_{F}^{2}}+\lambda_{z}\alpha_{z}\mathbf{K}_{grad}$ (13)
$\text{where }\textbf{K}_{grad}=\frac{2\mathbf{h}_{euc}^{\top}}{\left\|\mathbf{\sigma}_{euc}\right\|_{F}^{2}}\ast\left(\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}\right)$ (14)
2.5 Gradient descent step-size
Suppose a function $f:\mathbb{R}^{n\times n}\xrightarrow{}\mathbb{R}^{n\times n}$ which is convex and differentiable. Its gradient is Lipschitz continuous if there exists some constant C such that:
$\left\|\nabla f(\mathbf{x^{\prime}})-\nabla f(\mathbf{x})\right\|$ $\leq C\left\|\mathbf{x^{\prime}}-\mathbf{x}\right\|$
Since the loss functions 6-8 are convex and differentiable, one could find Lipschitz constants $C_{\{r,i,z\}}$ such that:

### v2

# Joint multiband deconvolution for Euclid and Vera C. Rubin images

@@
Joint multiband deconvolution for Euclid and Vera C. Rubin images
@@
With the advent of surveys like Euclid and Vera C. Rubin, astrophysicists will have access to both deep, high-resolution images and multiband images. However, these two types are not simultaneously available in any single dataset. It is therefore vital to devise image deconvolution algorithms that exploit the best of both worlds and that can jointly analyze datasets spanning a range of resolutions and wavelengths. In this work we introduce a novel multiband deconvolution technique aimed at improving the resolution of ground-based astronomical images by leveraging higher-resolution space-based observations. The method capitalizes on the fortunate fact that the Rubin $r$, $i$, and $z$ bands lie within the Euclid VIS band. The algorithm jointly de-convolves all the data to convert the $r$-, $i$-, and $z$-band Rubin images to the resolution of Euclid by leveraging the correlations between the different bands. We also investigate the performance of deep-learning-based denoising with DRUNet to further improve the results. We illustrate the effectiveness of our method in terms of resolution and morphology recovery, flux preservation, and generalization to different noise levels. This approach extends beyond the specific Euclid-Rubin combination, offering a versatile solution to improving the resolution of ground-based images in multiple photometric bands by jointly using any space-based images with overlapping filters.
@@
Deconvolution – Euclid – Vera C. Rubin – HST
@@ 1 Introduction
High spatial resolution, a high signal-to-noise ratio (S/N), and broad wavelength coverage are all essential for most observations in astrophysics. However, it is difficult, or even impossible, to have all three simultaneously. Space telescopes, although free of atmospheric turbulence, are limited in size. Ground-based telescopes can deliver high-S/N data but are affected by atmospheric turbulence and have a higher sky background. In addition, blurring by the instrumental or atmospheric point spread function (PSF) differs for each type of data and varies from band to band. To capitalize on the strengths of all types of telescopes and data, it is crucial to develop robust deconvolution techniques that can remove blurring by the PSF and optimize the S/N of the final reconstruction, by combining all observations and accounting for the different bands and PSFs.
@@
Due to the presence of noise, image deconvolution is a challenging ill-posed inverse problem that requires regularization for a well-defined solution. Early approaches in the field acknowledged this issue, proposing solutions such as minimizing the Tikhonov function (Tikhonov & Arsenin, 1977) or maximizing the entropy of the solution (Skilling & Bryan, 1984). Bayesian methods also emerged, including the Richardson-Lucy algorithm applied to early Hubble Space Telescope (HST) data (Richardson, 1972; Lucy, 1974). A novel approach proposed by Magain et al. (1998) separated point sources from extended ones and used a narrow PSF for deconvolution to achieve an improved resolution suitable for the chosen pixel sampling. This approach was improved with wavelet regularization for the extended channel (Cantale et al., 2016) and further refined by Michalewicz et al. (STARRED; 2023), who employed Starlets, an isotropic wavelet basis (Starck et al., 2015), to regularize the solution. There have also been efforts to jointly deconvolve multiple astronomical observations of the same sky region (Donath et al., 2023). Furthermore, Ingaramo et al. (2014) explored the combination of multiple sources by demonstrating the application of Richardson-Lucy deconvolution to merge high-resolution, high-noise images with low-resolution, low-noise images.
@@
A notable advancement in astronomical deconvolution was the use of deep learning. Once trained, deep-learning-based methods offer significant computational efficiency compared to traditional approaches. U-Nets (Ronneberger et al., 2015) have gained popularity for their nonlinear processing capabilities and multi-scale architecture. Expanding on U-Nets, Sureau et al. (2020) introduced the Tikhonet method for deconvolving optical galaxy images, demonstrating its superior performance over sparse regularization methods in terms of the mean squared error and a shape criterion that assesses galaxy ellipticity. Nammour et al. (2022) improved Tikhonet by incorporating a shape constraint into the loss function. Another powerful architecture, Learnlet (Ramzi et al., 2023), combines the strengths of wavelets and U-Nets while offering a fully interpretable neural network with minimal hallucination. In our previous work (Akhaury et al., 2022, 2024), we proposed a two-step deconvolution framework and investigated the performance of convolutional neural network (CNN) and transformer-based denoisers. We concluded that a Swin-transformer-based U-Net (SUNet; Fan et al. 2022) outperforms a CNN-based U-Net (Ronneberger et al., 2015) in terms of normalized mean squared error (NMSE) and structural similarity index measure.
@@
While deconvolution is primarily used to reconstruct galaxy images at high spatial resolution in each photometric band independently, there are scenarios, particularly at low S/Ns, where joint multiband deconvolution can enhance the detection and characterization of systems. One such potential application is the joint multiband deconvolution of Euclid and Vera C. Rubin images. The Rubin Observatory is set to deliver a dataset of $500$ petabytes across multiple optical frequency bands, while Euclid will observe images spanning the optical and infrared spectrum. Interestingly, the Euclid VIS band (central frequency $=715$ nm) overlaps with three of the Rubin filters: $r$, $i$, $z$. As a space-based satellite, Euclid will produce images with sharper details due to its narrower PSF compared to Rubin. A related study by Joseph et al. (2021) also involves jointly modeling simulated images that model observations from both Euclid and Rubin.
@@
In this work we present a novel multiband deconvolution technique designed to enhance the resolution of ground-based astronomical images by leveraging higher-resolution space-based observations. Our approach, which focuses on the joint deconvolution of Rubin and Euclid images, effectively exploits the overlapping spectral coverage of the Rubin $r$, $i$, and $z$ bands with the Euclid VIS band. By utilizing the Euclid VIS-band image as a term that provides additional information, our technique ensures that the deconvolved Rubin images retain high spatial resolution and accurate photometric measurements. The integration of deep-learning-based denoising further enhances the quality of the deconvolved outputs, reducing background noise without altering the main structures of the galaxies. We generated realistic Euclid and Rubin simulations from HST cutouts of varying magnitudes extracted from the GOODS-N and GOODS-S surveys (Retzlaff, J. et al., 2010). The simulated Euclid-like VIS-band PSF was obtained from Liaudat et al. (2022), and the simulated Rubin-like $r$-, $i$-, and $z$-band PSFs from Abolfathi et al. (2021). Our method is effective in terms of resolution recovery, flux preservation, and generalization across different noise levels. Through our joint deconvolution approach, we achieve resolution recovery in simulated Rubin images close to that of HST, a feat nearly impossible with independent deconvolutions of each photometric band. The potential applications of our method go beyond the Euclid-Rubin pair, providing a flexible solution to enhancing the resolution of ground-based images in multiple photometric bands with any overlapping space-based filter band. This versatility is especially important as large-scale astronomical surveys gather increasing amounts of data, creating a need for effective and reliable deconvolution techniques.
@@
In Sect. 2 we describe the deconvolution problem and introduce our proposed solution. The methodology for generating our dataset is detailed in Sect. 3. We then present the outcomes of our deconvolution algorithm in Sect. 4. Finally, Sect. 5 presents our conclusions. To support reproducible research, the codes utilized in this article are publicly accessible.
@@ 2.1 The forward model
For the three Rubin filters, let $\mathbf{y}_{r},\mathbf{y}_{i}$, and $\mathbf{y}_{z}\in\mathbb{R}^{n\times n}$ be the corresponding observed images and $\mathbf{h}_{r},\mathbf{h}_{i}$, and $\mathbf{h}_{z}\in\mathbb{R}^{n\times n}$ be the PSFs. If $\mathbf{x}_{r}^{t},\mathbf{x}_{i}^{t}$, and $\mathbf{x}_{z}^{t}\in\mathbb{R}^{n\times n}$ denote the corresponding target images, $\ast$ denotes the convolution operation, and $\eta_{r},\eta_{i}$, and $\eta_{z}\in\mathbb{R}^{n\times n}$ denote additive noise, the observed Rubin images can then be modeled as
@@
$\mathbf{y}_{z}$ $=\mathbf{h}_{z}\ast\mathbf{x}_{z}^{t}+\eta_{z}.$ (3)
As for Euclid, let $\mathbf{y}_{euc}\in\mathbb{R}^{n\times n}$ be the observed image, $\mathbf{x}^{t}_{euc}\in\mathbb{R}^{n\times n}$ be the target image, and $\mathbf{h}_{euc}\in\mathbb{R}^{n\times n}$ be the PSF. If $\eta_{euc}\in\mathbb{R}^{n\times n}$ denotes additive noise and $\alpha_{r},\alpha_{i},\alpha_{z}\in\mathbb{R}^{n}$ denote the corresponding fractional flux contribution from each Rubin filter, the target and the observed images can be modeled as
@@
$\mathbf{y}_{euc}$ $=\mathbf{h}_{euc}\ast\mathbf{x}_{euc}^{t}+\eta_{euc.}$ (5)
The motivation behind taking the weighted sum of the Rubin images to model the Euclid image can be seen in Fig. 1, which shows the overlap between the Rubin and Euclid filters.
Figure 1: Filter curves for Euclid and Rubin.  The relative filter transmission is shown as a function of the wavelength. The Euclid VIS band overlaps with three of the Rubin filters: $r$, $i$, and $z$.
@@ 2.2 The proposed solution
$L_{z}(\mathbf{x}_{z})$ $=\frac{1}{2}\left\|\frac{\mathbf{h}_{z}\ast\mathbf{x}_{z}-\mathbf{y}_{z}}{\sigma_{z}}\right\|_{F}^{2}+\lambda_{z}\textbf{K,}$, (8)
where $\text{where }\textbf{K}=\left\|\frac{\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}}{\sigma_{euc}}\right\|_{F}^{2}.$ (9)
The first terms in Eqs. 6-8 represent the data fidelity terms for each respective band, with $\sigma_{r}$, $\sigma_{i}$, and $\sigma_{z}$ being the corresponding noise maps. The second terms are the constraining terms that enforce the condition that the sum of the Rubin $r$-, $i$-, and $z$-band images equals the Euclid VIS-band image, as expected from Fig. 1 after flux calibration. Within these constraining terms, the individual images $\mathbf{x}_{r}$, $\mathbf{x}_{i}$, and $\mathbf{x}_{z}$ are multiplied by their respective fractional flux contributions $\alpha_{r}$, $\alpha_{i}$, and $\alpha_{z}$, which represent the fractional area overlaps between their corresponding filter curves. The values of $\alpha_{r}$, $\alpha_{i}$, and $\alpha_{z}$ are obtained by integrating the area under the curves in Fig. 1 and normalizing them to sum up to one. The resulting values are $\alpha_{r}=0.3785$, $\alpha_{i}=0.3468$, and $\alpha_{z}=0.2746$. The denominator, $\sigma_{euc}$, denotes the Euclid image noise map. The choice of the multiplicative hyperparameters $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ is described in Sect. 2.3.
@@ 2.3 Hyper-parameter tuning
The hyper-parameters $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ are determined by varying the ratios between the second and first terms in Eqs. 6-8. These ratios indicate the contribution coming from the constraining term. We varied the ratios from 0 (no contribution from the constraining term) to 1 (equal contribution from the constraining term) in steps of 0.01. From this experiment, the optimal solution yields the lowest mean squared error with a ratio of 0.3 for all three photometric bands. Subsequently, the values of $\lambda_{r}$, $\lambda_{i}$, and $\lambda_{z}$ are computed from these ratios by dividing them by K (Eq. 9).
@@ 2.4 Optimization
The aim is to find optimal solutions $\mathbf{\hat{x}}_{r}$, $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{x}}_{z}$ that minimize the individual loss functions:
$\mathbf{\hat{x}}_{\{r,i,z\}}=\underset{\mathbf{x}_{\{r,i,z\}}}{\operatorname{argmin}}\hskip 2.0ptL_{\{r,i,z\}}(\mathbf{x}_{\{r,i,z\}}),$
@@
$\mathbf{x}_{\{r,i,z\}}^{[k+1]}=\mathbf{x}_{\{r,i,z\}}^{[k]}-\beta_{\{r,i,z\}}\nabla L_{\{r,i,z\}}(\mathbf{x}_{\{r,i,z\}}^{[k]}),$ (10)
where $\mathbf{x}^{[k]}$ denotes the variable $\mathbf{x}$ at $k^{th}$ iteration, and $\beta_{r},\beta_{i},\beta_{z}\in\mathbb{R}^{n}$ are the step sizes chosen such that convergence is guaranteed (described in more detail in Sect. 2.5). While computing Eq. 10 for one band, it is assumed that the other two bands are known and remain constant. The gradients of the loss functions 6-8 are given by
@@
$\nabla L_{z}(\mathbf{x}_{z})$ $=\frac{\mathbf{h}_{z}^{\top}\ast(\mathbf{h}_{z}\ast\mathbf{x}_{z}-\mathbf{y}_{z})}{\left\|\mathbf{\sigma}_{z}\right\|_{F}^{2}}+\lambda_{z}\alpha_{z}\mathbf{K}_{grad},$ (13)
$\text{where }\textbf{K}_{grad}=\frac{2\mathbf{h}_{euc}^{\top}}{\left\|\mathbf{\sigma}_{euc}\right\|_{F}^{2}}\ast\left(\hskip 2.0pt\mathbf{h}_{euc}\ast\sum\limits_{c\in\{r,i,z\}}\alpha_{c}\mathbf{x}_{c}-\mathbf{y}_{euc}\right).$ (14)
2.5 Gradient descent step size
Suppose a function $f:\mathbb{R}^{n\times n}\xrightarrow{}\mathbb{R}^{n\times n}$ that is convex and differentiable. Its gradient is Lipschitz-continuous if there exists some constant C such that
$\left\|\nabla f(\mathbf{x^{\prime}})-\nabla f(\mathbf{x})\right\|$ $\leq C\left\|\mathbf{x^{\prime}}-\mathbf{x}\right\|.$
Since the loss functions 6-8 are convex and differentiable, one could find Lipschitz constants $C_{\{r,i,z\}}$ such that